OpenStudy (anonymous):

solve for X 2(3x-5)=3(4x+1)

OpenStudy (anonymous):

distribute the 2 through, then distribute the 3 through

OpenStudy (anonymous):

Then subtract an x term from both sides.

OpenStudy (anonymous):

6x - 10 = 12x +3 12x - 6x =-10 -3 6x=-13

OpenStudy (anonymous):

x=-13/6

OpenStudy (anonymous):

thnks